# The effect of accretion on scalar superradiant instability

### v1

@@ III.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
The important quantities in the numerical evolution can be estimated with simple analytic expressions at $\alpha\ll 1$ limit. To facilitate the analysis, we also solve the evolution equations without GW emission. The results are shown with orange dashed curves in Fig. 3. The exponential rise of the BH mass in Eq. (19) is also displayed as a thin blue line in Fig. 3 (d). For compactness, the quantities at a specific time are labeled with the corresponding subscripts, with subscript 0 reserved for the initial time, e.g. $\alpha_{0}\equiv M_{0}\mu$, $M_{1}\equiv M(t_{1})$ and $M_{\text{s,0}^{\prime}}\equiv M_{\text{s}}(t_{0}^{\prime})$.
@@
$t_{0}^{\prime\prime}=\tau_{\mathrm{acc}}\log\left(\sqrt{6}k\right).$ (36)
In the range from $t=0$ to $t_{0}^{\prime\prime}$, the BH mass grows according to Eq. (19) until $M=\sqrt{6}kM_{0}$, and the BH spin follows Eq. (13) up to $a_{*}\approx a_{*\lim}$. Ignoring GW emission and inserting Eq. (25) and Eq. (13) into Eq. (11c), one could calculate the condensate mass
$\begin{split}M^{(011)}_{\text{s}}=M^{(011)}_{\text{s0}}\exp\left\{{\alpha_{0}^{8}\mu\tau_{\text{acc}}}\left[g\left(\frac{M}{M_{0}}\right)-g\left(1\right)\right]\right\},\hskip 14.22636pt&\\ \text{for }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k},&\end{split}$ (37)
@@
$\begin{split}g(x)\equiv&\ \frac{1}{36}\bigg{[}-\frac{2\alpha_{0}x^{9}}{3}+\frac{2}{7}\sqrt{6}kx^{7}\\ &+\frac{\sqrt{3}k}{35}\left(216k^{4}+36k^{2}x^{2}+5x^{4}\right)(9k^{2}-x^{2})^{3/2}\bigg{]}.\end{split}$ (38)
@@
Subsequently, the BH spin remains at approximately $a_{*\lim}$ from $t_{0}^{\prime\prime}$ to $t_{0}^{\prime}$. As the condensate mass increases, the BH spin evolution is increasingly affected by superradiance, which results in $\dot{a}_{*}=0$ at $t_{0^{\prime}}$. In this time range, one could simplify $\Gamma_{011}$ in Eq. (25) by taking the leading order (LO) of the $\alpha$ and setting $a_{*}=1$. Then Eq. (11c) can be solved by ignoring the GW emission, arriving at
$\begin{split}M^{(011)}_{\text{s}}&=M_{\text{s,0}^{\prime\prime}}\exp\left[\frac{\mu\,\tau_{\text{acc}}}{192}\left(\alpha^{8}-\alpha_{0^{\prime\prime}}^{8}\right)\right],\end{split}$ (39)
where $\alpha_{0^{\prime\prime}}=\sqrt{6}k\alpha_{0}$ and $M_{\text{s,0}^{\prime\prime}}$ is determined by Eq. (37).
@@
The evolution of the BH spin $a_{*}$ can be derived from Eqs. (11a) and (11b)
$\dot{a}_{*}=\frac{1}{\tau_{\text{acc}}}\left(\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}\right)-\frac{2M_{\text{s}}}{M}\left(\frac{1}{M\omega_{011}}-2a_{*}\right)\Gamma_{011},$ (40)
where the photon capture is ignored. From $t_{0}^{\prime\prime}$ to $t_{0}^{\prime}$, both terms on the right side are small, resulting in an almost constant $a_{*}$. From $t_{0}^{\prime}$ to $t_{1}$, the accretion term is still small, while the superradiance term grows quickly, leading to the fast drop of the BH spin. Thus one could ignore the accretion term in the time range between $t_{0}^{\prime\prime}$ and $t_{1}$. The BH mass $M$ can be estimated with Eq. (19) and the condensate mass continues to evolve according to the form given by Eq. (39). Then, using $\Gamma_{011}$ at the LO of the $\alpha$ and setting $a_{*}=1$, one rewrites Eq. (40) as
$\begin{split}\dot{a}_{*}=-&\frac{M_{\mathrm{s,0^{\prime\prime}}}^{(011)}}{24}M^{6}\mu^{8}\exp\left[\frac{\tau_{\text{acc}}}{192}(M^{8}-M^{8}_{0^{\prime\prime}})\mu^{9}\right].\end{split}$ (41)
Integrating both sides of Eq. (41) from $t_{0^{\prime\prime}}$, one obtains the following solution
$\begin{split}a_{*}(\alpha)&=a_{*0^{\prime\prime}}-\frac{1}{192}M^{(011)}_{\mathrm{s0^{\prime\prime}}}\tau_{\mathrm{acc}}\mu^{2}e^{-\frac{\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}}\\ &\,\times\left[\alpha_{0^{\prime\prime}}^{6}E_{\frac{1}{4}}\left(\frac{-\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)-\alpha^{6}E_{\frac{1}{4}}\left(\frac{-\alpha^{8}\tau_{\mathrm{acc}}\mu}{192}\right)\right],\end{split}$ (42)
where $E_{\frac{1}{4}}(x)$ denotes the exponential integral function. At $t_{1}$, the BH spin decreases to a value close to $a_{\mathrm{*c}}^{(011)}\sim 4\alpha_{1}\ll a_{*\lim}$. Thus, one can approximate $a_{*1}-a_{*0^{\prime\prime}}$ by $-a_{*\lim}$, and perform the following expansion
$\begin{split}E_{\frac{1}{4}}\left(\frac{-\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)&=\\ e^{\frac{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}{192}}&\left\{-\frac{192}{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}+\mathcal{O}\left[\left(\frac{1}{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}\right)^{2}\right]\right\}.\end{split}$ (43)
@@
$\alpha_{1}=\left\{\frac{-48}{\tau_{\text{acc}}\mu}W_{-1}\left[-\left(\frac{32\cdot 3^{3/4}\mu M_{\mathrm{s,0^{\prime\prime}}}^{(011)}\left(\tau_{\mathrm{acc}}\mu\right)^{1/4}}{\alpha_{0^{\prime\prime}}^{6}M_{\mathrm{s,0^{\prime\prime}}}^{(011)}\tau_{\mathrm{acc}}\mu^{2}E_{\frac{1}{4}}\left(\frac{-\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)-192a_{*0^{\prime\prime}}e^{\frac{\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}}}\right)^{4}\right]\right\}^{1/8},$ (44)
where $W_{-1}(x)$ is the Lambert $W$ function. If $M_{\text{s},0^{\prime\prime}}^{(011)}\mu\ll\alpha_{0^{\prime\prime}}^{2}$, the term with the Exponential integral function is a small contribution and can be dropped. With the value of $\alpha_{1}$, the time $t_{1}$ can be obtained using Eq. (19) and the condensate mass is estimated with Eq. (39).
The attractor phase spans from $t_{1}$ to $t_{3}$, when the BH mass reaches the critical value $M_{\text{de}}^{(011)}$ defined in Eq. (34). In this time range, the BH follows the Regge trajectory closely. We further separate this phase into two ranges with $t_{2}$. The condensate mass to BH mass ratio first increases with time, reaches the maximum at $t_{2}$, and then drops gradually due to the GW emission.
From $t_{1}$ to $t_{2}$, the photon absorption from the accretion disk is unimportant since the BH spin is smaller than $0.97$. The GW radiation has a small effect on $M_{s}/M$ at $t_{2}$, which can be seen from the difference of the black and red dashed curves at $t_{2}$ in Fig. 3. Below we treat this effect as perturbation. At the zeroth order, the GW radiation is ignored and the BH spin evolves along the Regge trajectory given by Eq. (10). Inserting Eqs. (11a) and (11b) into the derivative of $a_{*\text{c}}^{(011)}$ with respect to $t$, one could then solve the superradiance flux order by order in $\alpha=M\mu$
$\begin{split}\dot{E}_{\mathrm{SR}}^{(011)}=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\Bigg{(}3\sqrt{\frac{3}{2}}-\frac{31\alpha}{2}+\frac{787\alpha^{2}}{8\sqrt{6}}-\frac{20305\alpha^{3}}{216}&\\ +\frac{1564549\alpha^{4}}{3456\sqrt{6}}+\mathcal{O}(\alpha^{5})&\Bigg{)}.\end{split}$ (45)
Ignoring GW emission and using Eq. (11a), one could integrate this equation directly
$\begin{split}{{M}_{\mathrm{s,Regge}}^{(011)}(\alpha)}\equiv\frac{\alpha^{2}}{\mu}\Bigg{(}\frac{3}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{3}-\frac{473\alpha^{2}}{32\sqrt{6}}-\frac{223\alpha^{3}}{270}&\\ +\frac{999421\alpha^{4}}{20736\sqrt{6}}+\mathcal{O}(\alpha^{5})&\Bigg{)}.\end{split}$ (46)
@@
${{M}_{\mathrm{s,2}}^{(011),\text{LO}}}={M}_{\mathrm{s,1}}^{(011)}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})-{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{1}).$ (47)
@@
[…]
where Eqs. (19) and (46) have been used. Then the condensate mass at $t_{2}$ is
$\begin{split}{{M}_{\mathrm{s,2}}^{(011)}}&={{M}_{\mathrm{s,2}}^{(011),\text{LO}}}+\delta{{M}_{\mathrm{s,2}}^{(011)}}.\end{split}$ (49)
@@
$\left(1+\frac{{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}{M_{2}}\right)\dot{E}_{\mathrm{SR}}^{(011)}-\dot{E}_{\mathrm{GW}}^{(011)}=\frac{{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}{{\tau}_{\mathrm{acc}}}.$ (50)
Here we have assumed the condensate obtains most of its mass in the attractor phase, i.e., $M_{\text{s}}^{(011)}(\alpha_{2})\approx M_{\text{s,Regge}}^{(011)}(\alpha_{2})$. Inserting the expressions in Eqs. (45), (46) and (82) and assuming $\alpha_{2}\gg\alpha_{1}$, one finds the perturbation is invalid unless $\mu\tau_{\text{acc}}\sim\alpha_{2}^{14}$. Then the LO value of $\alpha_{2}$ is
$\alpha_{2}\approx 1.25\times\sqrt[14]{\frac{1}{\mu{\tau}_{\mathrm{acc}}}}.$ (51)
The value of $t_{2}$ can be estimated with Eq. (11a). Ignoring the photon-capturing and using the expression in Eq. (45), one arrives at
$\dot{M}\approx\frac{M}{{\tau}_{\mathrm{acc}}}-\dot{E}_{\mathrm{SR}}^{(011)}\approx\frac{M}{{\tau}_{\mathrm{acc}}}-3\sqrt{\frac{3}{2}}\frac{M^{2}\mu}{{\tau}_{\mathrm{acc}}},$ (52)
@@
$M\approx\frac{2M_{1}e^{(t-t_{1})/\tau}}{2+3\sqrt{6}M_{1}\mu\left[e^{(t-t_{1})/\tau}-1\right]}.$ (53)
@@
$t_{2}\approx t_{1}+{\tau}_{\mathrm{acc}}\log\left[\frac{\alpha_{2}(2-3\sqrt{6}\,\alpha_{1})}{\alpha_{1}(2-3\sqrt{6}\,\alpha_{2})}\right].$ (54)
@@
The second part of the attractor phase starts at $t_{2}$ and ends at $t_{3}$. In this time range, GW emission begins to dominate the evolution of the condensate. According to Eq. (82), the GW emission flux is determined by both the condensate mass ${M}_{\mathrm{s}}^{(011)}$ and the BH mass $M$. From $t_{2}$ to $t_{3}$, the former decreases over time while the latter grows almost exponentially. The resulting GW flux could increase, decrease or even be a constant, depending on the initial parameters. For the parameters used in Fig. 3, the GW emission flux is slightly enhanced in this time range. By definition, the corresponding BH mass $M_{3}\approx M_{\text{de}}^{(011)}$. The LO expression of $M_{\text{de}}^{(011)}$ is given in Eq. (34). Higher-order contributions of $M_{\text{de}}^{(011)}$ could be solved from
${a}_{\mathrm{*lim}}=\frac{4M_{3}\omega_{011}}{1+4M_{3}^{2}\omega_{011}^{2}},$ (55)
where $\omega_{011}$ can be found in Eq. (5) with $\alpha$ replaced by $\alpha_{3}$. The obtained mass coupling is $\alpha_{3}=M_{3}\mu\approx 0.469$. At $t_{3}$, Eq. (52) is not a good estimate due to the ignorance of the higher order terms. According to Eqs. (11a) and (11c), if GW emission and photon capturing are ignored, the sum of the BH mass and the condensate mass increases almost exponentially. Then $t_{3}$ can be estimated by
$t_{3}$ $=t_{2}+{\tau}_{\mathrm{acc}}\log\left[\frac{M_{3}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{3})}{M_{2}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}\right].$ (56)
@@
Num. $3.78\times 10^{-2}$ $3.35\times 10^{-2}$ $6.16\times 10^{7}$ $1.49\times 10^{-1}$ $9.75\times 10^{-2}$ $1.26\times 10^{8}$ $2.41\times 10^{8}$
$10^{-3}$ Est. $7.43\times 10^{-3}$ $1.25\times 10^{-2}$ $1.12\times 10^{10}$ $9.55\times 10^{-2}$ $5.94\times 10^{-2}$ $9.91\times 10^{10}$ $2.18\times 10^{11}$

### v2

@@ 3.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
The important quantities in the numerical evolution can be estimated with simple analytic expressions at $\alpha\ll 1$ limit. To facilitate the analysis, we also solve the evolution equations without GW emission. The results are shown with orange dashed curves in Fig. 3. The exponential rise of the BH mass in Eq. (2.19) is also displayed as a thin blue line in Fig. 3 (d). For compactness, the quantities at a specific time are labeled with the corresponding subscripts, with subscript 0 reserved for the initial time, e.g. $\alpha_{0}\equiv M_{0}\mu$, $M_{1}\equiv M(t_{1})$ and $M_{\text{s,0}^{\prime}}\equiv M_{\text{s}}(t_{0}^{\prime})$.
@@
$t_{0}^{\prime\prime}=\tau_{\mathrm{acc}}\log\left(\sqrt{6}k\right).$ (3.3)
In the range from $t=0$ to $t_{0}^{\prime\prime}$, the BH mass grows according to Eq. (2.19) until $M=\sqrt{6}kM_{0}$, and the BH spin follows Eq. (2.13) up to $a_{*}\approx a_{*\lim}$. Ignoring GW emission and inserting Eq. (2.25) and Eq. (2.13) into Eq. (2.11c), one could calculate the condensate mass
$M^{(011)}_{\text{s}}=M^{(011)}_{\text{s0}}\exp\left\{{\alpha_{0}^{8}\mu\tau_{\text{acc}}}\left[g\left(\frac{M}{M_{0}}\right)-g\left(1\right)\right]\right\},\text{ for }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k},$ (3.4)
@@
$g(x)\equiv$ $\ \frac{1}{36}\left[-\frac{2\alpha_{0}x^{9}}{3}+\frac{2}{7}\sqrt{6}kx^{7}+\frac{\sqrt{3}k}{35}\left(216k^{4}+36k^{2}x^{2}+5x^{4}\right)(9k^{2}-x^{2})^{3/2}\right].$ (3.5)
@@
Subsequently, the BH spin remains at approximately $a_{*\lim}$ from $t_{0}^{\prime\prime}$ to $t_{0}^{\prime}$. As the condensate mass increases, the BH spin evolution is increasingly affected by superradiance, which results in $\dot{a}_{*}=0$ at $t_{0^{\prime}}$. In this time range, one could simplify $\Gamma_{011}$ in Eq. (2.25) by taking the leading order (LO) of the $\alpha$ and setting $a_{*}=1$. Then Eq. (2.11c) can be solved by ignoring the GW emission, arriving at
$M^{(011)}_{\text{s}}=M_{\text{s,0}^{\prime\prime}}\exp\left[\frac{\mu\,\tau_{\text{acc}}}{192}\left(\alpha^{8}-\alpha_{0^{\prime\prime}}^{8}\right)\right],$ (3.6)
where $\alpha_{0^{\prime\prime}}=\sqrt{6}k\alpha_{0}$ and $M_{\text{s,0}^{\prime\prime}}$ is determined by Eq. (3.4).
@@
The evolution of the BH spin $a_{*}$ can be derived from Eqs. (2.11a) and (2.11b)
$\dot{a}_{*}=\frac{1}{\tau_{\text{acc}}}\left(\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}\right)-\frac{2M_{\text{s}}}{M}\left(\frac{1}{M\omega_{011}}-2a_{*}\right)\Gamma_{011},$ (3.7)
where the photon capture is ignored. From $t_{0}^{\prime\prime}$ to $t_{0}^{\prime}$, both terms on the right side are small, resulting in an almost constant $a_{*}$. From $t_{0}^{\prime}$ to $t_{1}$, the accretion term is still small, while the superradiance term grows quickly, leading to the fast drop of the BH spin. Thus one could ignore the accretion term in the time range between $t_{0}^{\prime\prime}$ and $t_{1}$. The BH mass $M$ can be estimated with Eq. (2.19) and the condensate mass continues to evolve according to the form given by Eq. (3.6). Then, using $\Gamma_{011}$ at the LO of the $\alpha$ and setting $a_{*}=1$, one rewrites Eq. (3.7) as
$\dot{a}_{*}=-\frac{M_{\mathrm{s,0^{\prime\prime}}}^{(011)}}{24}M^{6}\mu^{8}\exp\left[\frac{\tau_{\text{acc}}}{192}(M^{8}-M^{8}_{0^{\prime\prime}})\mu^{9}\right].$ (3.8)
Integrating both sides of Eq. (3.8) from $t_{0^{\prime\prime}}$, one obtains the following solution
$\begin{split}a_{*}(\alpha)=\,&a_{*0^{\prime\prime}}-\frac{1}{192}M^{(011)}_{\mathrm{s0^{\prime\prime}}}\tau_{\mathrm{acc}}\mu^{2}e^{-\frac{\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}}\\ &\times\left[\alpha_{0^{\prime\prime}}^{6}E_{\frac{1}{4}}\left(\frac{-\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)-\alpha^{6}E_{\frac{1}{4}}\left(\frac{-\alpha^{8}\tau_{\mathrm{acc}}\mu}{192}\right)\right],\end{split}$ (3.9)
where $E_{\frac{1}{4}}(x)$ denotes the exponential integral function. At $t_{1}$, the BH spin decreases to a value close to $a_{\mathrm{*c}}^{(011)}\sim 4\alpha_{1}\ll a_{*\lim}$. Thus, one can approximate $a_{*1}-a_{*0^{\prime\prime}}$ by $a_{*\lim}$, and perform the following expansion
$E_{\frac{1}{4}}\left(\frac{-\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)=e^{\frac{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}{192}}$ $\left\{-\frac{192}{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}+\mathcal{O}\left[\left(\frac{1}{\alpha_{1}^{8}\tau_{\mathrm{acc}}\mu}\right)^{2}\right]\right\}.$ (3.10)
@@
$\alpha_{1}=\left\{\frac{-48}{\tau_{\text{acc}}\mu}W_{-1}\left[-\left(\frac{32\cdot 3^{3/4}\mu M_{\mathrm{s,0^{\prime\prime}}}^{(011)}\left(\tau_{\mathrm{acc}}\mu\right)^{1/4}}{\alpha_{0^{\prime\prime}}^{6}M_{\mathrm{s,0^{\prime\prime}}}^{(011)}\tau_{\mathrm{acc}}\mu^{2}E_{\frac{1}{4}}\left(\frac{-\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}\right)-192a_{*0^{\prime\prime}}e^{\frac{\alpha_{0^{\prime\prime}}^{8}\tau_{\mathrm{acc}}\mu}{192}}}\right)^{4}\right]\right\}^{1/8},$ (3.11)
where $W_{-1}(x)$ is the Lambert $W$ function. If $M_{\text{s},0^{\prime\prime}}^{(011)}\mu\ll\alpha_{0^{\prime\prime}}^{2}$, the term with the Exponential integral function is a small contribution and can be dropped. With the value of $\alpha_{1}$, the time $t_{1}$ can be obtained using Eq. (2.19) and the condensate mass is estimated with Eq. (3.6).
The attractor phase spans from $t_{1}$ to $t_{3}$, when the BH mass reaches the critical value $M_{\text{de}}^{(011)}$ defined in Eq. (3.1). In this time range, the BH follows the Regge trajectory closely. We further separate this phase into two ranges with $t_{2}$. The condensate mass to BH mass ratio first increases with time, reaches the maximum at $t_{2}$, and then drops gradually due to the GW emission.
From $t_{1}$ to $t_{2}$, the photon absorption from the accretion disk is unimportant since the BH spin is smaller than $0.97$. The GW radiation has a small effect on $M_{s}/M$ at $t_{2}$, which can be seen from the difference of the black and red dashed curves at $t_{2}$ in Fig. 3. Below we treat this effect as perturbation. At the zeroth order, the GW radiation is ignored and the BH spin evolves along the Regge trajectory given by Eq. (2.10). Inserting Eqs. (2.11a) and (2.11b) into the derivative of $a_{*\text{c}}^{(011)}$ with respect to $t$, one could then solve the superradiance flux order by order in $\alpha=M\mu$
$\dot{E}_{\mathrm{SR}}^{(011)}=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left(3\sqrt{\frac{3}{2}}-\frac{31\alpha}{2}+\frac{787\alpha^{2}}{8\sqrt{6}}-\frac{20305\alpha^{3}}{216}+\frac{1564549\alpha^{4}}{3456\sqrt{6}}+\mathcal{O}(\alpha^{5})\right).$ (3.12)
Ignoring GW emission and using Eq. (2.11a), one could integrate this equation directly
${{M}_{\mathrm{s,Regge}}^{(011)}(\alpha)}\equiv\frac{\alpha^{2}}{\mu}\left(\frac{3}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{3}-\frac{473\alpha^{2}}{32\sqrt{6}}-\frac{223\alpha^{3}}{270}+\frac{999421\alpha^{4}}{20736\sqrt{6}}+\mathcal{O}(\alpha^{5})\right).$ (3.13)
@@
${{M}_{\mathrm{s,2}}^{(011),\text{LO}}}={M}_{\mathrm{s,1}}^{(011)}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})-{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{1}).$ (3.14)
@@
$\delta{{M}_{\mathrm{s,2}}^{(011)}}$ $=-\int^{t_{2}}_{t_{1}}\dot{E}_{\mathrm{GW}}^{(011)}(t)dt\approx-5.24\times 10^{-3}{{\tau}_{\mathrm{acc}}}\alpha_{2}^{16},$ (3.15)
where Eqs. (2.19) and (3.13) have been used. Then the condensate mass at $t_{2}$ is
${{M}_{\mathrm{s,2}}^{(011)}}$ $={{M}_{\mathrm{s,2}}^{(011),\text{LO}}}+\delta{{M}_{\mathrm{s,2}}^{(011)}}.$ (3.16)
@@
$\left(1+\frac{{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}{M_{2}}\right)\dot{E}_{\mathrm{SR}}^{(011)}-\dot{E}_{\mathrm{GW}}^{(011)}=\frac{{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}{{\tau}_{\mathrm{acc}}}.$ (3.17)
Here we have assumed the condensate obtains most of its mass in the attractor phase, i.e., $M_{\text{s}}^{(011)}(\alpha_{2})\approx M_{\text{s,Regge}}^{(011)}(\alpha_{2})$. Inserting the expressions in Eqs. (3.12), (3.13) and (A.5) and assuming $\alpha_{2}\gg\alpha_{1}$, one finds the perturbation is invalid unless $\mu\tau_{\text{acc}}\sim\alpha_{2}^{14}$. Then the LO value of $\alpha_{2}$ is
$\alpha_{2}\approx 1.25\times\sqrt[14]{\frac{1}{\mu{\tau}_{\mathrm{acc}}}}.$ (3.18)
The value of $t_{2}$ can be estimated with Eq. (2.11a). Ignoring the photon-capturing and using the expression in Eq. (3.12), one arrives at
$\dot{M}\approx\frac{M}{{\tau}_{\mathrm{acc}}}-\dot{E}_{\mathrm{SR}}^{(011)}\approx\frac{M}{{\tau}_{\mathrm{acc}}}-3\sqrt{\frac{3}{2}}\frac{M^{2}\mu}{{\tau}_{\mathrm{acc}}},$ (3.19)
@@
$M\approx\frac{2M_{1}e^{(t-t_{1})/\tau}}{2+3\sqrt{6}M_{1}\mu\left[e^{(t-t_{1})/\tau}-1\right]}.$ (3.20)
@@
$t_{2}\approx t_{1}+{\tau}_{\mathrm{acc}}\log\left[\frac{\alpha_{2}(2-3\sqrt{6}\,\alpha_{1})}{\alpha_{1}(2-3\sqrt{6}\,\alpha_{2})}\right].$ (3.21)
@@
The second part of the attractor phase starts at $t_{2}$ and ends at $t_{3}$. In this time range, GW emission begins to dominate the evolution of the condensate. According to Eq. (A.5), the GW emission flux is determined by both the condensate mass ${M}_{\mathrm{s}}^{(011)}$ and the BH mass $M$. From $t_{2}$ to $t_{3}$, the former decreases over time while the latter grows almost exponentially. The resulting GW flux could increase, decrease or even be a constant, depending on the initial parameters. For the parameters used in Fig. 3, the GW emission flux is slightly enhanced in this time range. By definition, the corresponding BH mass $M_{3}\approx M_{\text{de}}^{(011)}$. The LO expression of $M_{\text{de}}^{(011)}$ is given in Eq. (3.1). Higher-order contributions of $M_{\text{de}}^{(011)}$ could be solved from
${a}_{\mathrm{*lim}}=\frac{4M_{3}\omega_{011}}{1+4M_{3}^{2}\omega_{011}^{2}},$ (3.22)
where $\omega_{011}$ can be found in Eq. (2.5) with $\alpha$ replaced by $\alpha_{3}$. The obtained mass coupling is $\alpha_{3}=M_{3}\mu\approx 0.469$. At $t_{3}$, Eq. (3.19) is not a good estimate due to the ignorance of the higher order terms. According to Eqs. (2.11a) and (2.11c), if GW emission and photon capturing are ignored, the sum of the BH mass and the condensate mass increases almost exponentially. Then $t_{3}$ can be estimated by
$t_{3}$ $=t_{2}+{\tau}_{\mathrm{acc}}\log\left[\frac{M_{3}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{3})}{M_{2}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2})}\right].$ (3.23)
@@
Num. $78\times 10^{-2} 3.3$$35\times 10^{-2}$ $6.16\times 10^{7}$ $1.49\times 10^{-1}$ $9.75\times 10^{-2}$ $1.26\times 10^{8}$ $2.41\times 10^{8}$
$10^{-3}$ Est. $7.43\times 10^{-3}$ $25\times 10^{-2} 1.1$$12\times 10^{10}$ $9.55\times 10^{-2}$ $5.94\times 10^{-2}$ $9.91\times 10^{10}$ $2.18\times 10^{11}$